I need help with this question, thank you in advance!!!
A study has shown that 20% of all college textbooks cost at least $70. It is known that the standard deviation of the prices of the college textbook is $9.50. Assuming that prices of all college textbooks follows a normal distribution; find the mean price of all college textbooks.

Found 2 solutions by ewatrrr, rothauserc:
Answer by ewatrrr(24785)


Hi
invNorm(.80) = .8416 Note: 80% cost < $70
.8416 = (70-m)/9.5
9.5(.8416) = 70 - m
m = 70 - 9.5(.8416) = 70 - 7.995 = $62

Answer by rothauserc(4718)

We are given Pr(X<70) = 0.20 or 20%
consult z tables for z value associated with 0.20
z value is -0.84
we know that the z value is (X - mean of x) / standard deviation
therefore
-0.84 = (70 - mean) / 9.50
70 - mean = -0.84 * 9.50 = -7.98
-mean = -7.98 -70
mean = $77.98